OpenStudy (anonymous):

|3x + 2| = 14 My answer is x=4 and x=-4 Am I right?

OpenStudy (anonymous):

4 is correct -4 is not

OpenStudy (anonymous):

|3x + 2| = 14 3x + 2 = 14 or 3x + 2 = -14 3x = 12 x = 4 try the other one

OpenStudy (anonymous):

That isn't one of my options? They are A.x = 4 and x = −4 B.x = 4 and x = -16/3 c. x =28/3 and x = 16/3

OpenStudy (anonymous):

i know 3x + 2 = -14 solve for x and you will get it

OpenStudy (anonymous):

-4 ?

OpenStudy (anonymous):

show me how?